Please g(x)=log(x+2), graph it, equation for the vertical asymptote,x intercept in (x,y) form
Also the same things for g(x)=-log(x)
Thank you

Answer by stanbon(75887)

Please g(x)=log(x+2), graph it, equation for the vertical asymptote,x intercept in (x,y) form
Verticle asymptote: x = -2
x-intercept:
Let y = 0
log(x+2)=0
x+2 = 10^0
x+2 = 1
x = -1
x-intercept is (-1,0)
---------------------------
Also the same things for g(x)=-log(x)
Verticle asymptote: x=0
x-intercept
Let y=0
-logx = 0
logx = 0
x = 10^0
x = 1
x-intercept is (1,0)
=========================
Cheers,
Stan H.